pH of barium hydroxide solution is 12. It`s solubility product is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From,pH + pOH = 14we get, pOH = 14 -12 = 2Now from formuale: pOH = -log[OH-][OH]Ba(OH)2---> Ba+2+ 2OH-when the reaction reaches at equilibrium pt.:Ba+2= x and OH-= 2xsince 2x =10-2as eq. 1therefore x = 10-2/2 = 0.5 * 10-2Hence ksp= [Ba+2]*[OH-]2= [0.5* 10-2][10-2]2= 0.5 *10-6=5 *10-7 (approx.)